### v1

@@ 7 Conclusion and Discussion
Tuberculosis (TB), a highly contagious airborne disease, was declared a global public health emergency by the WHO in 1993. Since then, numerous global initiatives have been undertaken to combat and prevent TB. The WHO’s End TB Strategy aims to reduce the burden of the TB epidemic in alignment with the United Nations’ Sustainable Development Goal 3: Good health and well-being. Despite these efforts, the 2024 Global TB Report highlighted a worrying increase in new TB cases, rising globally from 7.5 million in 2022 to 8.2 million in 2023. While low and middle-income countries remain high-burden regions, the impacts of rapid globalization have contributed to an uptick in TB cases in high-income countries. Japan, with the world’s highest aging population and growing migration from TB-endemic regions, has shifted to a medium-burden country with rising TB incidence, presenting a significant public health and economic burden. Early detection and forecasting systems are vital for effective TB control. This study introduces a technological solution combining the networked SIR model with advanced deep learning-based forecasting architectures. These frameworks provide actionable insights for forecasting TB cases in Japan. Our method bridges the gap between compartmental models, which describe disease dynamics based on core epidemiological drivers, and data-driven forecasting techniques, which predict future disease trajectories. Integrating the networked SIR model with data-driven deep learning frameworks allows for precise TB incidence forecasts while capturing the interactions of key disease drivers, spatial dynamics, and historical patterns. This ensures that our models provide realistic forecasts for designing effective public health interventions and making real-time policy adjustments during an outbreak.
@@
The networked SIR model, incorporating a saturated incidence rate and graph Laplacian diffusion, offers a robust estimation of TB transmission across Japan’s 47 prefectures. The model’s positivity and boundedness properties ensure realistic epidemic projections, while the global stability analysis reinforces its applicability to various populations and epidemic scenarios. Experimental results demonstrate that the EGDL-based models consistently outperform traditional data-driven models, yielding accurate forecasts across different time horizons (for short to medium-term forecasts). Moreover, the probabilistic outputs of the EGDL models provide valuable insights, allowing public health officials to track disease trends and tailor intervention strategies. Our approach provides a comprehensive understanding of future TB dynamics, excelling in short to medium-term forecasting, while the networked SIR model offers reliable long-term trajectory estimations. By forecasting TB incidence across Japan’s 47 prefectures, this model identifies potential hotspots, guiding the design of region-specific interventions such as vaccination campaigns, resource allocation, and public health policy adjustments. While this study focuses on TB in Japan, the proposed frameworks can be adapted for other countries with varying epidemiological profiles and population dynamics. Additionally, the models can be extended to address the dynamics of vector-borne or waterborne diseases. Future research could also explore the application of these frameworks to multi-disease contexts, such as TB-HIV and TB-diabetes co-infections, from both global and regional perspectives. Future investigations would examine how factors like socioeconomic status, healthcare access, and environmental variables influence epidemic forecasts, improving the precision of intervention strategies.
@@ Appendix A Mathematical Proofs
    for $u\geq 0,\,a>c$. By applying the Lyapunov-LaSalle invariance principle [50], we obtain ${\lim_{t\to\infty}u(t)=u^{*}=\dfrac{a-c}{\alpha c}}$.
@@
Since $S(x,0)$ and $I(x,0)$ in $\mathbb{D}_{+}$, Eq. (25) yields that there are no solutions $\{(S,I):x\in{V}\}$ that leave the hyperplanes from $S=0,I=0$. The vector field is either pointing to the interior of $\mathbb{D}_{+}$ or tangent to the hyperplane. Therefore, all the solutions starting in $\mathbb{D}_{+}$ stay in $\mathbb{D}_{+}$ for all time $t>0$, i.e., the region $\mathbb{D}_{+}$ is positively invariant.
@@ A.5 Proof of the Theorem 1
For any $\epsilon>0$ satisfying Eq. (27), substituting Eq. (29) into equation of (11b) and for $t>T_{1}$, we deduce
@@ A.6 Proof of the Theorem 2
$B_{1}^{S}=\dfrac{1}{\mu}\left[\Lambda-\dfrac{\beta A_{1}^{S}A_{1}^{I}}{1+\alpha A_{1}^{I}}\right]$ (42)
@@
$\limsup_{t\to\infty}S(x,t)\leq\dfrac{1}{\mu}\left[\Lambda-\dfrac{\beta(B_{1}^{S}-\epsilon)(B_{1}^{I}-\epsilon)}{1+\alpha(B_{1}^{I}-\epsilon)}\right],\quad\mbox{ uniformly in }x\in V.$
@@
A few simple computations and the condition $\mathcal{R}_{0}>1$ lead $B_{1}^{S}>0$. Thus, for any $\epsilon>0$ there is $T_{7}>T_{6}$ such that if $t>T_{7}$, $S(x,t)>B_{2}^{S}-\epsilon$, uniformly in $x\in V$. Again Eq. (11b)), when $t>T_{7},$ it follows
@@
If we carrying-on similar arguments, we acquire four sequences $\{A_{n}^{S}\}_{n=1}^{\infty},\{A_{n}^{I}\}_{n=1}^{\infty},\{B_{n}^{S}\}_{n=1}^{\infty}$ and $\{B_{n}^{I}\}_{n=1}^{\infty}$, where
@@ Appendix B Workflow of EGDL Frameworks
The workflow of the EGDL-Parallel and EGDL-Series frameworks are presented in Algorithm 1 and Fig. A1 presents a schematic representation of these architectures.
Input : Incidence cases recorded at $n$ locations in $T$ consecutive timestamps as $Y(x,t);\;x=1,2,\ldots,n;\;t=1,2,\ldots,T$.
Output : $q$-step ahead forecast of the disease incidence cases for the $n$ locations as $\left\{\widehat{Y}(1,T+v),\widehat{Y}(2,T+v),\ldots,\widehat{Y}(n,T+v)\right\};\;v=1,2,\ldots,q$ ($\geq 1$).
@@
1 Build a networked SIR epidemic model with saturated incidence rate and graph Laplacian diffusion by considering key disease drivers across $n$ locations.
2 Calculate the posterior distribution of the networked SIR epidemic model parameters using the MCMC approach and select the best-fitted parameters to quantify the spatial dynamics of disease transmission and the effect of key epidemic drivers.
3 Solve the networked SIR model to estimate the infected curve across $n$ locations for $\tau$ timestamps as $\left\{I(1,\tau),I(2,\tau),\ldots,I(n,\tau)\right\};\;\tau=1,2,\ldots,T+q$.
@@
2 Provide a hybrid input to $f_{EGDLP}$ framework combining the historical incidence cases $Y(x,t)$ and the estimated infected curve of networked SIR $I(x,t)$, formulated as $\left\{\left\{Y\left(x,\tilde{t}-t_{0}\right)\right\}_{t_{0}=0}^{t_{w}-1},\;\left\{I\left(x,\tilde{t}-t_{0}\right)\right\}_{t_{0}=0}^{t_{w}-1}\right\}_{\tilde{t}=t_{w}}^{T},$ with $t_{w}$ lagged time steps, where $Y(x,t)$ serves as endogenous variable and $I(x,t)$ acts as the exogenous variable.
@@
1 Obtain the residuals of the networked SIR model by subtracting the estimated infected curve from the ground truth observations as ${e(x,t)=Y(x,t)-I(x,t)};\;x=1,2,\ldots,n,\;t=1,2,\ldots,T$.
32 Initialize a data-driven forecasting model $f_{EGDLS}\in\mathcal{F}$ and provide the $t_{w}$ lagged values of the residual series as input to generate the q-step ahead forecasts as:
$\hat{e}(x,\tilde{t}+v)=f_{EGDLS}\left(\left\{e\left(x,\tilde{t}-t_{0}\right)\right\}_{t_{0}=0}^{t_{w}-1}\right),\;\tilde{t}=t_{w},\ldots,T;\;v=1,2,\ldots,q.$
Combine the deterministic predictions of the networked SIR model $I(x,\tilde{t}+v)$, with the forecasts of stochastic components from $f_{EGDLS}$ framework as
$\widehat{Y}(x,\tilde{t}+v)=I(x,\tilde{t}+v)+\hat{e}(x,\tilde{t}+v),$
to obtain the $q$-step ahead forecast of the disease incidence cases.
@@
Figure A1: Epidemic-Guided Deep Learning (EGDL). The upper panel of the plot represents a schematic architecture of the networked SIR model, and the lower panel showcases the workflow of the EGDL-Parallel framework (lower left) and the EGDL-Series approach (lower right).
@@ Appendix C Empirical Results
Table A4 summarizes the descriptive statistics of Japan TB data, including measures of central tendency, standard deviation (Sd), coefficient of variation (CV), skewness, and kurtosis. Additionally, to investigate key statistical characteristics of these datasets, we study their global patterns, as outlined below: Long-term dependency reflects the self-similarity of the time series with its lagged observations. This feature is crucial for probabilistic time series modeling approaches. In this study, the Hurst exponent is computed to detect whether TB incidence trends exhibit long-range dependence over time. Stationarity ensures the mean and variance of a time series remain constant over time, implying no systematic changes in trend or seasonal patterns. This property is essential for many classical forecasting models. The Kwiatkowski–Phillips–Schmidt–Shin (KPSS) test is employed to assess the level and trend stationarity of the TB incidence data. Linearity of a time series determines whether its future values can be expressed as a linear combination of past observations. This feature influences model selection based on their ability to capture linear or non-linear dynamics. The Teräsvirta neural network test is used in this study to detect the presence of non-linear patterns in TB incidence cases. Seasonality represents recurring patterns within a defined time interval. We apply the Ollech and Webel combined seasonality test to identify seasonal behavior in TB incidence. All the statistical analyses were performed using R statistical software, following [36].
(a) Laplacian matrix (b) Correlation Plot
Refer to caption Refer to caption
Figure A2: (a) The Laplacian matrix corresponding to the network depicted in Fig. 1(A), representing the connectivity and structure of the nodes in the system. This matrix captures the relationships between neighboring nodes and is essential for analyzing diffusion processes and other network dynamics. (b) Pairwise correlation of TB incidence cases recorded at 47 prefectures of Japan.
@@
Station No. Descriptive Statistics Statistical Properties
@@ C.1 Parameter Calibration of the Mechanistic Model
Fig. A3 depicts the stability plot of the disease-free equilibrium ($\mathcal{R}_{0}=0.38$) and endemic equilibrium ($\mathcal{R}_{0}=3.8$). The time evolution of the infected individuals for selected prefectures of Japan estimated using the modified networked SIR model is presented in Fig. A4.
@@
C.2 Causality Analysis
To assess temporal predictive causality, we apply the Granger causality (GC) test, which determines whether the lagged values of one variable enhance the explanatory power of another variable beyond what is provided by its own lagged values [58]. This linear test evaluates the null hypothesis of Granger causation and is particularly suited for identifying predictive relationships. Since the GC test requires time series data to be stationary and linear, conditions that are frequently violated in the case of Japan’s TB data (Table A4), we also utilize wavelet coherence analysis [59]. This non-parametric method identifies dynamic, time-dependent interactions between two signals in the time-frequency domain. It is especially effective for detecting localized correlations and co-movements in non-stationary time series. The p-values obtained from the GC test range between $\left(4.77\times 10^{-20},1.00\times 10^{-3}\right)$, with most values falling below $2.00\times 10^{-4}$ across different prefectures in Japan. Since these p-values are below the significance level of 0.05, the null hypothesis is rejected, indicating Granger’s causation between the networked SIR model’s infection curve and the TB incidence cases. Additionally, Fig. A5 illustrates the wavelet coherence across significant time-frequency domain pairs for selected prefectures. In these plots, the horizontal axis represents the time (in months), while the vertical axis represents the frequency dimension. Warmer colors highlight regions with significant co-movement between the two time series, while cooler colors indicate weaker dependencies. This visualization reveals both the frequency and timing of significant interactions.
Figure A5: Wavelet coherence of TB incidence cases and the infected curve predicted by the networked SIR model for selected prefectures of Japan. The horizontal axis depicts the monthly time, while the vertical axis shows the period. The warmer the color of a region, the higher the degree of dependence between the pair.
C.3 Performance Measures
The mathematical expression for calculating the error metrics for the $x^{th}$ prefecture are as follows:

### v2

@@ 8 Conclusion and Discussion
This study introduces a technological solution combining the MN-SIR model with advanced deep learning-based forecasting architectures to predict TB incidence cases. Our method bridges the gap between compartmental models, which leverage disease dynamics based on core epidemiological drivers, and data-driven forecasting techniques, which predict future disease trajectories. Integrating the MN-SIR model with deep learning frameworks allows for precise TB incidence forecasts while capturing the interactions of key disease drivers, spatial dynamics, and historical patterns. This ensures that our models provide accurate and realistic forecasts for designing effective public health interventions and making real-time policy adjustments during an outbreak.
@@
The MN-SIR model, incorporating a saturated incidence rate and graph Laplacian diffusion, offers a robust estimation of TB transmission across Japan’s 47 prefectures. The model’s positivity and boundedness properties ensure realistic epidemic projections, while the global stability analysis reinforces its applicability to various populations and epidemic scenarios. Experimental results demonstrate that the EGDL-based models consistently outperform traditional data-driven models, yielding accurate forecasts across different time horizons (for short to medium-term forecasts). Moreover, the probabilistic outputs of the EGDL models, along with forecast explainability achieved through temporal Grad-CAM, provide valuable and transparent insights that help public health officials to track disease trends and tailor intervention strategies. Our approach provides a comprehensive understanding of future TB dynamics, excelling in short to medium-term forecasting, while the MN-SIR model offers reliable long-term trajectory estimations. By forecasting TB incidence across Japan’s 47 prefectures, this model identifies potential hotspots, guiding the design of region-specific interventions such as vaccination campaigns, resource allocation, and public health policy adjustments. To support the generalizability of our framework, we conducted additional experiments on TB incidence data from mainland China. The EGDL models maintained superior forecasting performance, demonstrating robustness to different population dynamics and surveillance contexts. These cross-country results confirm the broader applicability of our framework and validate the title’s claim of epidemic-guided forecasting for TB outbreaks in diverse settings. Additionally, the models can be extended to address the dynamics of vector-borne or waterborne diseases. Future research could also explore the application of these frameworks to multi-disease contexts, such as TB-HIV and TB-diabetes co-infections, from both global and regional perspectives. Further investigations may also examine how factors like socio-economic status, healthcare access, and environmental variables influence epidemic forecasts, improving the precision of intervention strategies.
Data and Code Availability Statement
Acknowledgments
We are grateful to the editor, guest editors, and the three anonymous reviewers for their thoughtful comments and constructive feedback, which greatly improved the quality of this work. T.C. would also like to thank his student, Ms. Donia Besher, for her valuable assistance in collecting the data from China.
@@ Appendix A Mathematical Proofs
    for $u\geq 0,\,a>c$. We construct a suitable Lyapunov function $W(u(t))$ that is continuously differentiable and positive definite with respect to the equilibrium $u^{*}$. We then show that its time derivative along the trajectories of the system satisfies ${W}^{\prime}(u(t))\leq 0$, indicating that $W(u(t))$ is non-increasing over time. Hence, by applying the Lyapunov-LaSalle invariance principle [55], we obtain ${\lim_{t\to\infty}u(t)=u^{*}=\dfrac{a-c}{\alpha c}}$.
@@
The term $\sum_{\begin{subarray}{c}y\in V\\ y\sim x\end{subarray}}(S(y,t)-S(x,t))\geq 0$, using Eq. (5); here the negative part becomes zero as $S(x,t)=0$. A similar argument can be made for $I$. Since $S(x,0)$ and $I(x,0)$ in $\mathbb{D}_{+}$, Eq. (25) yields that there are no solutions $\{(S,I):x\in{V}\}$ that leave the hyperplanes from $S=0,I=0$. The vector field is either pointing to the interior of $\mathbb{D}_{+}$ or tangent to the hyperplane. Therefore, all the solutions starting in $\mathbb{D}_{+}$ stay in $\mathbb{D}_{+}$ for all time $t>0$, i.e., the region $\mathbb{D}_{+}$ is positively invariant.
@@ A.6 Proof of the Theorem 2
$B_{1}^{S}=\dfrac{1}{\mu}\left[\Lambda-\dfrac{\beta A_{1}^{S}A_{1}^{I}}{1+\alpha A_{1}^{I}}\right].$ (42)
@@
$\limsup_{t\to\infty}S(x,t)\leq\dfrac{1}{\mu}\left[\Lambda-\dfrac{\beta(B_{1}^{S}-\epsilon)(B_{1}^{I}-\epsilon)}{1+\alpha(B_{1}^{I}-\epsilon)}\right],\quad\mbox{ uniformly in }x\in V$
@@
A few simple computations and the condition $\mathcal{R}_{0}>1$ lead $B_{1}^{S}>0$. Thus, for any $\epsilon>0$ there is $T_{7}>T_{6}$ such that if $t>T_{7}$, $S(x,t)>B_{2}^{S}-\epsilon$, uniformly in $x\in V$. Again Eq. (12b), when $t>T_{7},$ it follows
@@
If we carry on similar arguments, we acquire four sequences $\{A_{n}^{S}\}_{n=1}^{\infty},\{A_{n}^{I}\}_{n=1}^{\infty},\{B_{n}^{S}\}_{n=1}^{\infty}$ and $\{B_{n}^{I}\}_{n=1}^{\infty}$, where
@@ Appendix B Workflow of EGDL Frameworks
The workflow of the EGDL-Parallel and EGDL-Series frameworks is presented in Algorithm 1.
Input : Incidence cases recorded at $n$ locations in $T$ consecutive timestamps as $\left\{Y(x,t)\right\}_{x\in V,\;t\in\mathcal{T}};\;V=\{1,2,\ldots,n\};\;\mathcal{T}=\{1,2,\ldots,T\}$.
Output : $q$-step ahead forecast of the disease incidence cases for the $n$ locations as $\left\{\widehat{Y}(x,T+v)\right\}_{x\in V};\;v=1,2,\ldots,q$ ($\geq 1$).
@@
1 Build a MN-SIR epidemic model with saturated incidence rate and graph Laplacian diffusion by considering disease drivers across $n$ locations.
2 Calculate the posterior distribution of the MN-SIR model parameters using the MCMC approach and select the best-fitted parameters to quantify the spatial dynamics of disease transmission and the effect of key epidemic drivers.
3 Solve the MN-SIR model to estimate the infected curve across $n$ locations for $\tilde{\tau}$ timestamps as $\left\{I(x,\tilde{\tau})\right\}_{x\in V};\;\tilde{\tau}=1,2,\ldots,T+q$.
@@
2 Provide a hybrid input to $f_{EGDLP}$ framework combining the historical incidence cases $Y(x,t)$ and the estimated infected curve of MN-SIR $I(x,t)$, formulated as $\left\{\left\{Y\left(x,\tilde{t}-t_{0}\right)\right\}_{t_{0}=0}^{t_{w}-1},\;\left\{I\left(x,\tilde{t}-t_{0}\right)\right\}_{t_{0}=0}^{t_{w}-1}\right\}_{\tilde{t}=t_{w}}^{T},$ with $t_{w}$ lagged time steps, where $Y(x,t)$ serves as endogenous variable and $I(x,t)$ acts as the exogenous variable.
@@
1 Obtain the residuals of the MN-SIR model by subtracting the estimated infected curve from the ground truth observations as ${e(x,t)=Y(x,t)-I(x,t)};\;x\in V,\;t\in\mathcal{T}$.
32 Initialize a data-driven forecasting model $f_{M}\in\mathcal{F}$ and provide the $t_{w}$ lagged values of the residual series as input to generate the q-step ahead forecasts as:
$\hat{e}(x,\tilde{t}+v)=f_{M}\left(\left\{e\left(x,\tilde{t}-t_{0}\right)\right\}_{t_{0}=0}^{t_{w}-1}\right),\;\tilde{t}=t_{w},\ldots,T;\;v=1,2,\ldots,q.$
Combine the deterministic components of the MN-SIR model $I(x,\tilde{t}+v)$ with the forecasts of stochastic components to generate the predictions of the EGDL-Series methods as:
$\widehat{Y}(x,\tilde{t}+v)=\hat{f}_{EGDLS}(x,\tilde{t}+v)=I(x,\tilde{t}+v)+\hat{e}(x,\tilde{t}+v).$
@@ Appendix C Empirical Results
Tables A8 and A9 summarize the descriptive statistics, including measures of central tendency (Mean), standard deviation (Sd), coefficient of variation (CV), skewness, and kurtosis for TB datasets of Japan and China, respectively. Additionally, to investigate key statistical characteristics of these datasets, we study their global patterns, as outlined below: Long-term dependency reflects the self-similarity of the time series with its lagged observations. This feature is crucial for probabilistic time series modeling approaches. In this study, the Hurst exponent is computed to detect whether TB incidence trends exhibit long-range dependence over time. Stationarity ensures the mean and variance of a time series remain constant over time, implying no systematic changes in trend or seasonal patterns. This property is essential for many classical forecasting models. The Kwiatkowski–Phillips–Schmidt–Shin (KPSS) test is employed to assess the level and trend stationarity of the TB incidence data. Linearity of a time series determines whether its future values can be expressed as a linear combination of past observations. This feature influences model selection based on their ability to capture linear or non-linear dynamics. The Teräsvirta neural network test is used in this study to detect the presence of non-linear patterns in TB incidence cases. Seasonality represents recurring patterns within a defined time interval. We apply the Ollech and Webel combined seasonality test to identify seasonal behavior in TB incidence. All the statistical analyses were performed using R statistical software, following [41].
@@
Prefecture No. Descriptive Statistics Statistical Properties
@@
Table A9: Global features of TB incidence cases at 31 prefectures of China. In the table, (A) indicate annual seasonality.
Province No. Descriptive Statistics Statistical Properties
Mean Range Sd CV Skewness Kurtosis Hurst Exponent Characteristics
1 583.23 (436, 706) 69.13 0.12 -0.18 -0.96 0.60 Stationary, Nonlinear, Seasonality (A)
2 263.13 (171, 344) 38.28 0.15 -0.57 0.13 0.55 Stationary, Linear, Seasonality (A)
3 2798.79 (2155, 3369) 291.60 0.10 -0.16 -0.52 0.67 Stationary, Linear, Seasonality (A)
4 1252.46 (857, 1625) 183.02 0.15 -0.03 -0.44 0.68 Non-stationary, Linear
5 1015.60 (733, 1311) 148.08 0.15 0.23 -0.58 0.54 Stationary, Linear, Seasonality (A)
6 1975.31 (1357, 2584) 263.55 0.13 0.23 -0.20 0.58 Stationary, Linear, Seasonality (A)
7 1243.17 (737, 1732) 234.19 0.19 0.18 -0.48 0.66 Non-stationary, Linear, Seasonality (A)
8 2619.27 (1873, 3598) 437.06 0.17 0.17 -0.67 0.69 Non-stationary, Linear, Seasonality (A)
9 552.88 (435, 657) 65.29 0.12 -0.24 -1.03 0.50 Stationary, Linear, Seasonality (A)
10 2503.90 (1718, 3193) 333.38 0.13 -0.12 -0.38 0.72 Non-stationary, Linear, Seasonality (A)
11 2318.58 (1745, 2825) 302.19 0.13 -0.23 -1.11 0.62 Stationary, Linear, Seasonality (A)
12 2951.65 (2314, 3721) 340.14 0.12 0.18 -0.52 0.62 Non-stationary, Linear, Seasonality (A)
13 1413.56 (1046, 1741) 178.00 0.13 -0.23 -0.74 0.61 Non-stationary, Linear, Seasonality (A)
14 3194.77 (1976, 24882) 3218.01 1.01 6.79 46.67 0.51 Stationary, Linear, Seasonality (A)
15 2649.19 (1882, 4059) 443.85 0.17 0.88 1.38 0.62 Non-stationary, Linear, Seasonality (A)
16 4877.46 (3688, 5769) 470.66 0.10 -0.10 -0.04 0.68 Non-stationary, Linear, Seasonality (A)
17 3656.77 (2695, 4550) 460.29 0.13 -0.23 -0.50 0.63 Non-stationary, Linear, Seasonality (A)
18 4508.90 (3443, 5801) 587.06 0.13 0.24 -0.73 0.65 Non-stationary, Linear, Seasonality (A)
19 6687.92 (4987, 7976) 729.69 0.11 -0.38 -0.76 0.57 Stationary, Linear, Seasonality (A)
20 3691.96 (2726, 5086) 554.50 0.15 0.30 -0.38 0.64 Non-stationary, Linear, Seasonality (A)
21 689.88 (485, 857) 87.91 0.13 -0.27 -0.42 0.68 Non-stationary, Linear, Seasonality (A)
22 1892.35 (1328, 2592) 289.06 0.15 0.09 -0.23 0.58 Stationary, Linear, Seasonality (A)
23 4523.88 (3586, 6044) 535.51 0.12 0.47 0.32 0.55 Non-stationary, Linear, Seasonality (A)
24 3784.88 (2774, 5251) 546.27 0.14 0.33 0.47 0.53 Stationary, Linear, Seasonality (A)
25 2201.13 (1616, 2997) 298.40 0.14 0.34 0.71 0.54 Stationary, Linear, Seasonality (A)
26 399.88 (176, 553) 85.51 0.21 -0.47 0.18 0.59 Stationary, Linear, Seasonality (A)
27 1809.21 (1462, 2261) 202.04 0.11 0.30 -0.95 0.58 Stationary, Linear, Seasonality (A)
28 1265.38 (921, 2039) 238.73 0.19 0.88 0.94 0.57 Stationary, Linear, Seasonality (A)
29 595.15 (375, 831) 106.01 0.18 0.14 -0.64 0.66 Non-stationary, Linear, Seasonality (A)
30 243.67 (146, 394) 48.72 0.20 0.92 1.47 0.66 Non-stationary, Linear, Seasonality (A)
31 3638.77 (2212, 5142) 689.76 0.19 0.19 -0.73 0.58 Stationary, Linear, Seasonality (A)
@@ C.1 Parameter Calibration of the Mechanistic Model
Fig. A1 and A3 depicts the stability plot of the disease-free equilibrium ($\mathcal{R}_{0}=0.38$) and endemic equilibrium ($\mathcal{R}_{0}=3.8$) for Japan and China, respectively. The time evolution of the infected individuals for selected locations of Japan and China, estimated using the MN-SIR model is presented in Fig. A2 and A4, respectively.
@@
(a) Time evolution for $S$ (b) Time evolution for $I$
Refer to caption Refer to caption
(c) Time evolution for $S$ (d) Time evolution for $I$
Refer to caption Refer to caption
Figure A3: (a) - (b) illustrates the stability plot of the disease-free equilibrium for $\mathcal{R}_{0}=0.38$ ($<1$) with China network. (c) - (d) illustrates the stability plot of the endemic equilibrium for $\mathcal{R}_{0}=3.8$ ($>1$) with China network.
(a) Province 20 (b) Province 10
Refer to caption Refer to caption
Figure A4: Time evolution of infected individuals with China network at (a) province 20 and (b) province 10 when the saturation factor ($\alpha$) varies from $0.1$ to $1.0$.
C.2 Performance Measures
The mathematical expression for calculating the error metrics corresponding to location $x$ is defined as follows: